OpenStudy (anonymous):

I suck at these! If you could answer that would be great! : Which statement is equivalent to |x - 9| < 17?

OpenStudy (anonymous):

x - 9 < 17 x - 9 < -17 x - 9 < 17 or x - 9 > -17 -17 < x - 9 < 17

OpenStudy (anonymous):

sorry maybe not lemme check

OpenStudy (anonymous):

i dont think thats correct.. :/

OpenStudy (anonymous):

well im not sure.

OpenStudy (anonymous):

its coreect3

OpenStudy (anonymous):

alright. Thanks. :)

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

The correct answer is #4. The #3 is incorrect because it should be AND not OR. -17 < x - 9 < 17

OpenStudy (anonymous):

well sorry i read it in haste( actually in my language(hindi), or means and )

OpenStudy (anonymous):

u also wrote or:P:P

OpenStudy (anonymous):

Typo

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

If |x - 9| = 17, then x - 9 = -17 OR x - 9 =17 If |x - 9| < 17, then we can write in 2 ways: x -9 > -17 AND x - 9 < 17, or -17 < x - 9 < 17 If |x - 9| > 17, then x - 9 < -17 OR x - 9 > 17

OpenStudy (anonymous):

It has "Delete" but no "Edit" button. If we made one letter or number typo, we have to retype the entire solution!

OpenStudy (anonymous):

or else copy paste nd edit